OpenStudy (anonymous):

1. How are the functions y = x and y = x – 3 related? How are their graphs related? (1 point)

OpenStudy (anonymous):

Each output for y = x – 3 is 3 more than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units. Each output for y = x – 3 is 3 more than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated up 3 units. Each output for y = x – 3 is 3 less than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units. Each output for y = x – 3 is 3 less than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated up 3 units.

OpenStudy (anonymous):

x-3 is shifted down by 3. So they are parallel.

OpenStudy (anonymous):

So, Each output for y = x – 3 is 3 less than the corresponding output for y = x. The graph of y = x – 3 is the graph of y = x translated down 3 units.

OpenStudy (anonymous):

thanks!!! can you help me with some more?

OpenStudy (anonymous):

@q12157